∫20x(5x2+x-2) dx

∫4(102x^2) dx


1
Expert's answer
2021-04-26T08:16:57-0400

"\int(100x^3+20x^2-40x)dx=25x^4+20x^3/3-20x^3+C"


"4\smallint10^{2x^2}dx"

"u=\sqrt{2}x"

"du=\sqrt2dx"

"4\smallint10^{2x^2}dx=\frac{4}{\sqrt2}\smallint10^{u^2}du=4\frac{\sqrt{π}}{ 2^{3/2}\sqrt{ln(10)}}\intop\frac{2\sqrt{ln(10)}10^{u^2}}{\sqrtπ}du"

"\intop\frac{2\sqrt{ln(10)}10^{u^2}}{\sqrtπ}du=erfi(\sqrt{ln(10)}u)"

"4\smallint10^{2x^2}dx=4\frac{\sqrt{π}}{ 2^{3/2}\sqrt{ln(10)}}erfi(\sqrt{ln(10)}u)=4\frac{\sqrt{π}}{ 2^{3/2}\sqrt{ln(10)}}erfi(\sqrt{ln(10)}\sqrt2x)+C"